# Identification of a Kalman Filter: Consistency of Local Solutions

### v2

@@ 6 Extensions
where $\beta\in{\mathbb{R}}^{{n_{\beta}}}$ is an unknown parameter vector to be estimated, and $\Phi_{k}\in{\mathbb{R}}^{{n}\times{n_{\beta}}}$ are some known regression matrices. The parameters to be estimated are denoted by $\theta\coloneqq(\beta,L,S)$. Here, the predicted states also depend on $\beta$:
@@
The results from Section 4 still hold for the optimization problem (26), under some additional Persistent Excitation (PE) conditions on the regressors $\Phi_{k}$, and the assumption that they are independent of the noise $e_{k}$.
Unfortunately, the latter assumption excludes the case $\Phi_{k}=\Phi(u_{k})$ when the data come from a closed-loop experiment.
@@
The expected value of the cost function $J_{N}(\theta)$ of the optimization problem (26) can be expressed as a sum:
${\mathbb{E}}\left[J_{N}(\theta)\right]$ $=\bar{J}^{\beta}_{N}(\beta,L,S)+\bar{J}^{{\mathrm{L}}}_{N}(L,S)+\bar{J}^{{\mathrm{S}}}_{N}(S),$ (27)
where the functions $\bar{J}^{\beta}_{N}(\beta,L,S)$, $\bar{J}^{{\mathrm{L}}}_{N}(L,S)$, and $\bar{J}^{{\mathrm{S}}}_{N}(S)$ are expressed as:
$\bar{J}^{\beta}_{N}(\beta,L,S)$ $\coloneqq\left\lVert\beta-\beta^{\star}\right\rVert^{2}_{Q_{N}(L,S)},$ (28a)
$\bar{J}^{{\mathrm{L}}}_{N}(L,S)$ $\coloneqq\frac{1}{N}\sum_{k=1}^{N}{\mathbb{E}}\left[\left\lVert Cz_{k}(L)\right\rVert^{2}_{S^{-1}}\right],$ (28b)
@@
for some positive-definite matrix $Q_{N}(L,S)$ that depends on the regressors $\Phi_{k}$.
@@ 7 Conclusion
This paper established that, when estimating only the Kalman gain of a known linear system, the PEM objective becomes asymptotically unimodal, and therefore remains a reliable identification method even when only local optimality can be guaranteed. The numerical examples support this conclusion: spurious minimizers disappear as the dataset grows, and standard optimization routines consistently converge to the true gain. The sketched extension to jointly estimating additional linear parameters and noise covariances is also encouraging, and formal proofs for this case is a future research direction.
Several open questions remain, including the rate at which local optima approach the true parameters as the amount of data increases. Another promising direction is the development of optimization algorithms tailored to this problem class; for instance, alternating or sequential updates over model parameters, Kalman gains, and covariances may yield fast and guaranteed convergence.
Appendix A Proof of Proposition 3
Proof A.1
@@ Appendix B Uniform law of large numbers
Consider the family of sequences $\{s_{k}(\theta),\,\theta\in\Theta\}$ defined as follows:
$s_{k}(\theta)=\sum_{i=0}^{k}H_{k,i}(\theta)w_{k-i},$ (32)
where $w_{k}$ are independent random variables with bounded fourth-order moments, and $\{H_{k,i}(\theta),\,\theta\in\Theta\}$ is a family of uniformly stable filters. More precisely, there exist some constants $c_{H},c_{w}>0$ and $\lambda\in(0,1)$ such that for all $k,i\in{\mathbb{N}}$ and all $\theta\in\Theta$, we have:
${\mathbb{E}}\left[\left\lVert w_{k}\right\rVert^{4}\right]\leq c_{w},\quad\text{and}\quad\left\lVert H_{k,i}(\theta)\right\rVert\leq c_{H}\lambda^{i}.$ (33)
@@
Appendix C Proof of Lemma 4
Proof C.1
We will prove the convergence result (13) by first showing that $\nabla V_{N}(L)$ converges to its expected value, and then proving that this expected value converges to $\nabla\bar{V}(L)$.
It can easily be verified that $e_{k}-Cz_{k}(L)$ satisfies the conditions of Lemma 9 with $\theta=L$ and $\Theta={\mathcal{L}}$. For any pair of indices $(i,j)$, the same holds for $\frac{\partial z_{k}(L)}{\partial L_{ij}}$. Thus, applying this lemma leads to the first desired result:
@@ Appendix C Proof of Lemma 4
Next, we express the expected value of $V_{N}(L)$ as:
${\mathbb{E}}\left[V_{N}(L)\right]$ $=$ (36)
${\mathrm{Tr}}{\left(W{\left[S^{\star}+C{\left(\textstyle\frac{1}{N}\sum\limits_{k=1}^{N}{\mathrm{Cov}}\left[z_{k}(L)\right]\right)}C^{\!\top}\!\right]}\right)},$
and ${\mathrm{Cov}}\left[z_{k}(L)\right]$ satisfies:
@@
$\qquad+(L-L^{\star})S^{\star}(L-L^{\star})^{\!\top}\!.$
Comparing these equations with (12-11), we find that:
${\mathbb{E}}\left[V_{N}(L)\right]-\bar{V}(L)={\mathrm{Tr}}{\left(\!WC{\left(\textstyle\frac{1}{N}\sum\limits_{k=1}^{N}X_{k}(L)\right)}C^{\!\top}\!\right)},$ (38)
with $X_{k}(L)\coloneqq{\mathrm{Cov}}\left[z_{k}(L)\right]-\bar{\Sigma}(L)$ satisfying:
@@
From the uniform stability assumption (Assumption 2), one can prove that $X_{k}(L)$ and its derivatives uniformly converge to zero:
$\sup_{L\in{\mathcal{L}}}\left\lVert\frac{\partial X_{N}(L)}{\partial L_{ij}}\right\rVert\xrightarrow[N\to+\infty]{}0.$ (40)
This result, combined with (38) leads to:
@@
Finally, combining the uniform convergence equations (35) and (41) we find that $\frac{\partial V_{N}(L)}{\partial L_{ij}}$ almost surely and uniformly converges to $\frac{\partial\bar{V}(L)}{\partial L_{ij}}$. Since this is true for any pair of indices $(i,j)$, the uniform convergence of the whole gradient is also established, which concludes the proof.

### v3

@@ 6 Extensions
where $\beta\in{\mathbb{R}}^{{n_{\beta}}}$ is an unknown parameter vector to be estimated, and $\Phi_{k}\in{\mathbb{R}}^{{n}\times{n_{\beta}}}$ are some known regression matrices. The parameters to be estimated are $\theta\coloneqq(\beta,L,S)$. Here, the predicted states also depend on $\beta$:
@@
As in (9), this cost function can be rewritten as follows:
$J_{N}(\beta,L,S)=\frac{1}{N}$ $\sum_{k=1}^{N}\left\lVert e_{k}-Cz_{k}(L)+\Psi_{k}(L)(\beta^{\star}-\beta)\right\rVert^{2}_{S^{-1}}$
$\hskip 42.67912pt+\log{\mathrm{det}}{\left(S\right)},$ (27)
with $\Psi_{k}(L)\coloneqq C\sum_{i=1}^{k}(A-LC)^{i-1}\Phi_{k-i}$, and $z_{k}(L)$ defined as in (10). Here, to ensure consistency of the estimates, we also need some Persistent Excitation (PE) condition for $N$ large enough and some stabilizing gain $\bar{L}$:
$\frac{1}{N}\sum_{k=1}^{N}\Psi_{k}(\bar{L})^{\!\top}\!\Psi_{k}(\bar{L}){\,\succcurlyeq\,}cI{\,\succ\,}0.$ (28)
One can prove, using some linear algebra, that when this condition holds for some $\bar{L}$, it also holds for any other $L$.
Furthermore, to prove asymptotic unimodality, we also need to assume that the noise $e_{k}$ is independent of the regressors $\Phi_{k}$, which excludes the case $\Phi_{k}=\Phi(u_{k})$ when the data come from a closed-loop experiment unfortunately.
@@
The results from Section 4 still hold for the optimization problem (26), under the assumption given above.
@@
Using (6), the decomposition ${\mathbb{E}}\left[J_{N}(\theta)\right]=\bar{J}^{\beta}_{N}(\beta,L,S^{-1})+\bar{J}^{{\mathrm{L}}}_{N}(L,S^{-1})+\bar{J}^{{\mathrm{S}}}_{N}(S)$ holds true, with the following terms:
$\bar{J}^{\beta}_{N}(\beta,L,W)$ $\coloneqq\left\lVert\beta-\beta^{\star}\right\rVert^{2}_{Q_{N}(L,W)},$ (29a)
$\bar{J}^{{\mathrm{L}}}_{N}(L,W)$ $\coloneqq\frac{1}{N}\sum_{k=1}^{N}{\mathbb{E}}\left[\left\lVert Cz_{k}(L)\right\rVert^{2}_{W}\right],$ (29b)
@@
with $Q_{N}(L,W)\coloneqq\frac{1}{N}\sum_{k=1}^{N}\Psi_{k}(L)^{\!\top}\!W\Psi_{k}(L)$.
@@ 7 Conclusion
This paper established that, when estimating only the Kalman gain of a known linear system, the PEM objective becomes asymptotically unimodal, and therefore remains a reliable identification method even when only local optimality can be guaranteed. The numerical examples support this conclusion: spurious minimizers disappear as the dataset grows, and standard optimization routines consistently converge to the true gain. The sketched extension to jointly estimating additional linear parameters and noise covariances is also encouraging, and formal proofs for such extensions are a future research direction.
Several open questions remain, for example, regarding the influence of model mismatch or the convergence speed of the local optima. Another promising direction is the development of optimization algorithms tailored to this problem class; for instance, alternating or sequential updates over model parameters, Kalman gains, and covariances may yield fast and guaranteed convergence.
Appendix A Technical proofs
Proof A.1 (Proof of Proposition 3)
@@ Appendix A Technical proofs
Let $w_{k}$ be independent random variables with $\sup\limits_{k\in{\mathbb{N}}}\,{\mathbb{E}}\left[\left\lVert w_{k}\right\rVert^{4}\right]<\infty$}.
Let $H(\theta)$ be a family of uniformly stable filters over $\theta\in\Theta$, i.e., for some $\lambda\in(0,1)$ and $c_{H}>0$, $\left\lVert H_{i}(\theta)\right\rVert\leq c_{H}\lambda^{i}$} holds for all $i\in{\mathbb{N}}$ and $\theta\in\Theta$.
Also define $s_{k}(\theta)\coloneqq\sum_{i=0}^{k}H_{i}(\theta)w_{k-i}$.
@@
Proof A.2 (Proof of Lemma 4)
The idea of the proof is to first show $\nabla V_{N}(L)\to{\mathbb{E}}\left[\nabla V_{N}(L)\right]$, then we will show ${\mathbb{E}}\left[\nabla V_{N}(L)\right]\to\nabla\bar{V}(L)$, and conclude from this that the desired result $\nabla V_{N}(L)\to\nabla\bar{V}(L)$ as in (13) holds.
First, $s_{k}(\theta)\equiv e_{k}-Cz_{k}(L)$ satisfies the conditions of Lemma 9 with $\theta=L$ and $\Theta={\mathcal{L}}$. For any pair of indices $(i,j)$, the same holds for $\frac{\partial z_{k}(L)}{\partial L_{ij}}$. Thus, applying this lemma leads to the first desired result:
@@
Next, we express ${\mathbb{E}}\left[V_{N}(L)\right]$ and compare it with $\bar{V}$ in (12):
${\mathbb{E}}\left[V_{N}(L)\right]$ $={\mathrm{Tr}}{\left(W{\left[S^{\star}+C{\left(\textstyle\frac{1}{N}\sum\limits_{k=1}^{N}{\mathrm{Cov}}\left[z_{k}(L)\right]\right)}C^{\!\top}\!\right]}\right)}$
$=\bar{V}(L)+{\mathrm{Tr}}{\left(\!WC{\left(\textstyle\frac{1}{N}\sum\limits_{k=1}^{N}X_{k}(L)\right)}C^{\!\top}\!\right)},$ (35)
with $X_{k}(L)\coloneqq{\mathrm{Cov}}\left[z_{k}(L)\right]-\bar{\Sigma}(L)$. Using the the dynamics of $z_{k}(L)$ in (10) we can derive:
@@
$\hskip 56.9055pt+(L-L^{\star})S^{\star}(L-L^{\star})^{\!\top}\!.$
Comparing this with the recursion for $\bar{\Sigma}(L)$ in (11), we find that $X_{k}(L)$ satisfies:
@@
Using the uniform stability from Assumption 2, equation (37) implies that $X_{k}(L)$ converges to zero uniformly over $L\in{\mathcal{L}}$. This result, combined with (A.2) leads to:
@@
By combining this with the uniform convergence equations (34), it follows that $\frac{\partial V_{N}(L)}{\partial L_{ij}}\to\frac{\partial\bar{V}(L)}{\partial L_{ij}}$ almost surely and uniformly. This holds for any pair of indices $(i,j)$ so the uniform convergence of the whole gradient is established.